OpenStudy (anonymous):

Find the value of y when x equals -2 in this equation. -3x + 3y = -12 A. 3 B. 6 C. -3 D. -6

OpenStudy (anonymous):

put x=-2 -3(-2)+3y=-12 6+3y=-12 3y=-12-6 3y=-18 y=-18/3=-6

OpenStudy (usukidoll):